OpenStudy (anonymous):

A larger storage unit is 20 ft long,12 ft high,and 10 ft wide. How much more volume does this storage unit have than the other one?

OpenStudy (anonymous):

Multiply again and subtract fromm 600?

OpenStudy (binary3i):

again do the same volume=lenght*breath*hight =20*12*10=2400 privios one was 600 therefor extra volume=2400-600=1800 ft^3

OpenStudy (anonymous):

thanks:)

OpenStudy (anonymous):

ive 3 more to go. tag?

OpenStudy (binary3i):

ok

OpenStudy (anonymous):

tnx:)